OpenStudy (anonymous):

dy/dx=(cosx)e^(y+sinx) instructions say to 'use separation of variables to solve the initial value problem' answer is y=ln(2-e^(sinx)) I get to y=-ln(sinx(e^sinx))+c without trouble, but I'm not sure where the 2 comes from or how to get rid of that extra sinx.

OpenStudy (tkhunny):

Let's see... \(\dfrac{dy}{dx} = \cos(x)e^{y+\sin(x)} = e^{y}\cos(x)e^{\sin(x)}\) \(\int e^{-y}dy = \int \cos(x)e^{\sin(x)}dx\) \(-e^{-y} = e^{\sin(x)} + C\) That's odd. \(e^{-y} > 0\), so logarithms are not appropriate. \(e^{-y} = -e^{\sin(x)} + D = D - e^{\sin(x)}\) Now, we can introduce the logarithm \( -y = \ln \left( D - e^{\sin(x)} \right) \) \( y = -\ln \left( D - e^{\sin(x)} \right) \) Looks to me like you're struggling with three things: 1) Just what is it with that Arbitrary Constant? What does it do? How is it used? 2) Some algebra issues. 3) Maybe the trigonometric antiderivative? I think we just need the initial conditions in order to get D = 2.

OpenStudy (anonymous):

Okay, thanks...one more question, in the third step, where does the cosx go? Wouldn't the antiderivative of that be sinx, so it would stick around? Other than that, very helpful though. I definitely have some issues with algebra, I'm kind of filling in the gaps as I go :)

OpenStudy (tkhunny):

See point #3? \(\dfrac{d}{dx}e^{\sin(x)} = e^{\sin(x)}\cdot \cos(x)\) You'll have to get up to speed on your chain rule for derivatives and your substitution methods for integrals.

OpenStudy (anonymous):

Okay, that could possibly have been very idiotic XD I get it now, thanks again!

OpenStudy (tkhunny):

No worries. We're all learning. Be careful and thorough. Point #1 is also very important. You just added "+ c" on the end after you were done. That's no good. It appears at the end of any indefinite integral. After that, you have to keep track of it differently.

OpenStudy (anonymous):

So just making sure, I need to add C right after I integrate and then treat it as a variable?

OpenStudy (tkhunny):

Not a variable, but as something that is a little less arbitrary than it was before. You can see that I changed from "C" to "D". This wasn't really necessary, but it was clear to me that I was about to change the sign of "C". Rather than write "-C" I decide to write "+D". This was just to emphasize that it was a different constant. A very common tranformatino of a constant is this one. ln(y) = ln(x) + C Applying Logarithms \(y = e^{ln(x)+C} = e^{ln(x)}e^{C} = Ae^{ln(x)} = Ax\) We had "C", but after the logarithm, we know it is a different constant, but we really don't know what it is, so we just renamed it.

OpenStudy (anonymous):

Okay, cool. Thanks!